OpenStudy (anonymous):

The speed at witch a certain computer can access the internet is 2 megabytes per seconde. How fast is this rate in megabytes per hour?

ganeshie8 (ganeshie8):

convert seconds to hour :- 60x60 seconds = 1 hour ?

OpenStudy (anonymous):

@Candy13106

OpenStudy (anonymous):

@ganeshie8 came you come back now? :)

OpenStudy (candy13106):

ok and than wat

OpenStudy (anonymous):

can*

ganeshie8 (ganeshie8):

in 1 hour, you got 60x60 = 3600 seconds

ganeshie8 (ganeshie8):

every second you're transmitting 2 megabytes so every 3600 seconds how many megabytes you wud be transmitting ?

OpenStudy (candy13106):

yea i already knew that

OpenStudy (candy13106):

um not quiet sure for that question

OpenStudy (candy13106):

u there

ganeshie8 (ganeshie8):

think a bit, its easy when u think.. me explaining it wil complicate :)

OpenStudy (candy13106):

ok than ill try

ganeshie8 (ganeshie8):

good luck :D

OpenStudy (candy13106):

so u wud do 3600 divide by 2 right

OpenStudy (candy13106):

ok than ill try

ganeshie8 (ganeshie8):

nopes, you wud multiply 2 megabytes wid 3600

ganeshie8 (ganeshie8):

that gives you how many megabytes are transmitted per hour (3600 seconds)

OpenStudy (candy13106):

oh ok so the oppsite right

OpenStudy (candy13106):

7200 right

ganeshie8 (ganeshie8):

yes ! 7200 megabytes per hour

OpenStudy (candy13106):

so is that the answer '

ganeshie8 (ganeshie8):

\(\large 2 \frac{MB}{sec}\) \(\large 2 \frac{MB}{sec} \times \frac{60 \times 60 sec}{hour}\) \(\large 2 \frac{MB}{\cancel{sec}} \times \frac{60 \times 60 \cancel{sec}}{hour}\) \(\large \frac{2 \times 60 \times 60 MB }{hour}\) \(\large \frac{7200 MB }{hour}\)

ganeshie8 (ganeshie8):

yes thats the answer, you can also write it as :- 7.2 gigabytes per hour

OpenStudy (candy13106):

yea and u got ot cancle out the hours right

ganeshie8 (ganeshie8):

nopes. hour stays, sec got cancelled cuz its there top and bottom

OpenStudy (candy13106):

well my math teacher said that u got to cancle out everything that has 2 of

ganeshie8 (ganeshie8):

your math teacher is right ! sec has 2, but hour has oly one.. :)

OpenStudy (candy13106):

ok well i guess ur right wats next

OpenStudy (candy13106):

or there is nothing else to do right

ganeshie8 (ganeshie8):

yes we're done :)

OpenStudy (candy13106):

ok well thanks alot